# Group class operations and homological conditions

### v1

@@ 4. Virtually Gorenstein group algebras
We fix a commutative ring $k$ and consider groups $G$ for which the group algebra $kG$ is virtually Gorenstein, i.e. for which ${\tt GProj}(kG)^{\perp}={{}^{\perp}}{\tt GInj}(kG)$. This homological condition on $kG$ has certain interesting consequences regarding the (homotopy) properties of Gorenstein modules. As an indication of these properties, we note that if $kG$ is virtually Gorenstein, then: [^2] Properties (a), (b), (c) and (d) are actually valid over any (left) virtually Gorenstein ring $R$.
(a) (cf. [37, Proposition 3.16]) There is an equality ${\tt GProj}(kG)={\tt PGF}(kG)$; hence, [35, Theorem 4.9] implies that the cotorsion pair $\left({\tt GProj}(kG),{\tt GProj}(kG)^{\perp}\right)$ is complete.
(b) (cf. [11, Proposition 2.6]) Let $M$ be a $kG$-module. Any resolution of $M$ which is obtained by iterating special Gorenstein projective precovers remains acyclic by applying the functor $\mbox{Hom}_{kG}(\_\!\_,N)$ for any Gorenstein injective $kG$-module $N$. Dually, any coresolution of $M$ which is obtained by iterating special Gorenstein injective preenvelopes remains acyclic by applying the functor $\mbox{Hom}_{kG}(L,\_\!\_)$ for any Gorenstein projective $kG$-module $L$.
(c) (cf. [11, Proposition 2.2]) An acyclic complex of $kG$-modules remains acyclic by applying the functor $\mbox{Hom}_{kG}(\_\!\_,N)$ for any Gorenstein injective $kG$-module $N$ if and only if it remains acyclic by applying the functor $\mbox{Hom}_{kG}(L,\_\!\_)$ for any Gorenstein projective $kG$-module $L$.
(d) (cf. [37, Example 3.18(2)]) Let $\mathcal{C}$ be the class of those acyclic complexes of $kG$-modules that satisfy the equivalent conditions described in (c) above. Then, the relative derived category, which is obtained from the homotopy category of $kG$ by inverting those chain maps whose mapping cone is a $\mathcal{C}$-complex, is triangulated equivalent to certain homotopy categories of Gorenstein projective (resp. Gorenstein injective) $kG$-modules.
We consider the classes $\mathfrak{X}$ and $\mathfrak{Y}$ defined in Section 2 and let $\mathfrak{Z}=\mathfrak{Z}(k)$ be the class consisting of those groups $G\in\mathfrak{X}\cap\mathfrak{Y}$, for which $kG$ is virtually Gorenstein.
@@
Proof. Let $G$ be a $\mathfrak{Z}$-group and consider a subgroup $H\subseteq G$. Then, $G\in\mathfrak{X}\cap\mathfrak{Y}$ and the group algebra $kG$ is virtually Gorenstein. Since $\mathfrak{X}$ and $\mathfrak{Y}$ are subgroup-closed classes (cf. Theorem 2.1), it follows that $H\in\mathfrak{X}\cap\mathfrak{Y}$. It remains to show that $kH$ is virtually Gorenstein.
Consider a $kH$-module $M\in{\tt GProj}(kH)^{\perp}$. Since $G\in\mathfrak{Z}\subseteq\mathfrak{X}$ and $\mathfrak{X}$ is subgroup-closed, we conclude that $H$ is an $\mathfrak{X}$-group. It follows that restriction maps ${\tt GProj}(kG)$ into ${\tt GProj}(kH)$ (this is precisely property (4) in Section 2), so that the coinduced $kG$-module $\mbox{coind}_{H}^{G}M$ is contained in ${\tt GProj}(kG)^{\perp}$, i.e. in ${{}^{\perp}}{\tt GInj}(kG)$. Invoking Corollary 2.2(ii), we may now conclude that $\mbox{res}_{H}^{G}\mbox{coind}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$. Since $M$ is a direct summand of $\mbox{res}_{H}^{G}\mbox{coind}_{H}^{G}M$, it follows that $M\in{{}^{\perp}}{\tt GInj}(kH)$. We have therefore shown that ${\tt GProj}(kH)^{\perp}\subseteq{{}^{\perp}}{\tt GInj}(kH)$.
Conversely, assume that $M\in{{}^{\perp}}{\tt GInj}(kH)$. Since $G\in\mathfrak{Z}\subseteq\mathfrak{Y}$ and $\mathfrak{Y}$ is subgroup-closed, we may conclude that $H$ is a $\mathfrak{Y}$-group. Hence, restriction maps ${\tt GInj}(kG)$ into ${\tt GInj}(kH)$ (this is precisely property (4’) in Section 2), so that the induced $kG$-module $\mbox{ind}_{H}^{G}M$ is contained in ${{}^{\perp}}{\tt GInj}(kG)$, i.e. in ${\tt GProj}(kG)^{\perp}$. Corollary 2.2(i) then implies that $\mbox{res}_{H}^{G}\mbox{ind}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$. Since $M$ is a direct summand of $\mbox{res}_{H}^{G}\mbox{ind}_{H}^{G}M$, it follows that $M\in{\tt GProj}(kH)^{\perp}$. We have thus shown that ${{}^{\perp}}{\tt GInj}(kH)\subseteq{\tt GProj}(kH)^{\perp}$. $\Box$
@@
Proof. (i) We have to show that ${\scriptstyle{{\bf LH}}}\mathfrak{Z}=\mathfrak{Z}$, i.e. that ${\scriptstyle{{\bf LH}}}\mathfrak{Z}\subseteq\mathfrak{Z}$. To that end, we consider an ${\scriptstyle{{\bf LH}}}\mathfrak{Z}$-group $G$ and aim at proving that $G\in\mathfrak{Z}$. First of all, we note that
$G\in{\scriptstyle{{\bf LH}}}\mathfrak{Z}\subseteq{\scriptstyle{{\bf LH}}}(\mathfrak{X}\cap\mathfrak{Y})\subseteq{\scriptstyle{{\bf LH}}}\mathfrak{X}\cap{\scriptstyle{{\bf LH}}}\mathfrak{Y}=\mathfrak{X}\cap\mathfrak{Y}.$
For any $\mathfrak{Z}$-subgroup $H\subseteq G$ we denote by $\left(\mbox{res}_{H}^{G}\right)^{-1}{\tt GProj}(kH)^{\perp}$ the class of those $kG$-modules $M$ for which $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ and apply Corollary 3.2 (with $\mathfrak{X}_{0}=\mathfrak{Z}\subseteq\mathfrak{X}$ therein), in order to conclude that
@@
Since ${\tt GProj}(kH)^{\perp}={{}^{\perp}}{\tt GInj}(kH)$ for any $\mathfrak{Z}$-subgroup $H\subseteq G$, the above equalities show that ${\tt GProj}(kG)^{\perp}={{}^{\perp}}{\tt GInj}(kG)$. The group algebra $kG$ is therefore virtually Gorenstein and hence $G\in\mathfrak{Z}$.
(ii) In order to show that $\Phi_{inj}\mathfrak{Z}=\mathfrak{Z}$, i.e. that $\Phi_{inj}\mathfrak{Z}\subseteq\mathfrak{Z}$, we consider a $\Phi_{inj}\mathfrak{Z}$-group $G$ and aim at proving that $G\in\mathfrak{Z}$. We have
$G\in\Phi_{inj}\mathfrak{Z}\subseteq\Phi_{inj}(\mathfrak{X}\cap\mathfrak{Y})\subseteq\Phi_{inj}\mathfrak{X}\cap\Phi_{inj}\mathfrak{Y}\subseteq\Phi_{flat}\mathfrak{X}\cap\Phi_{inj}\mathfrak{Y}=\mathfrak{X}\cap\mathfrak{Y},$
where the third inclusion follows since $\Phi_{inj}\mathfrak{X}\subseteq\Phi_{flat}\mathfrak{X}$; cf. [22, Lemma 1.5]. The group $G$ is contained in $\Phi_{inj}\mathfrak{Z}\subseteq\Phi_{flat}\mathfrak{Z}$ and hence we can apply Corollaries 3.2 and 3.5 in exactly the same way as in (i) above, in order to show that ${\tt GProj}(kG)^{\perp}={{}^{\perp}}{\tt GInj}(kG)$, i.e. that the group algebra $kG$ is virtually Gorenstein. This completes the proof that $G\in\mathfrak{Z}$. $\Box$
Let $\mathfrak{W}=\mathfrak{W}(k)$ be the class consisting of those groups $G$, for which $kG$ is weakly Gorenstein regular; in other words, $G\in\mathfrak{W}$ if the invariant $\mbox{sfli}\,kG$ is finite. This class contains all groups $G$, for which the group algebra $kG$ is Gorenstein regular. We note that $\mathfrak{W}$ is subgroup-closed; indeed, for any subgroup $H$ of a group $G$ there is an inequality $\mbox{sfli}\,kH\leq\mbox{sfli}\,kG$. Hence, $\mathfrak{W}$ is non-empty if and only if it contains the trivial group, i.e. if and only if the coefficient ring $k$ is itself weakly Gorenstein regular. In that case, $\mathfrak{W}$ is known to contain the class $\mathfrak{F}$ of finite groups; see, for example, [33].
@@
Proof. We note the following general facts, regarding a group $G$: If $F_{*}$ is an acyclic complex of flat $kG$-modules, then the complex of abelian groups $M\otimes_{kG}F_{*}$ is acyclic for any $kG$-module $M$ of finite flat dimension; this follows by induction on $\mbox{fd}_{kG}M$. Similarly, if $I^{*}$ is an acyclic complex of injective $kG$-modules and $M$ is a $kG$-module of finite flat dimension, then the complex of abelian groups $\mbox{Hom}_{kG}(M,I^{*})$ is acyclic; cf. [34, Corollary 5.9].
In order to prove that $\mathfrak{W}\subseteq\mathfrak{Z}$, we have to show that any $\mathfrak{W}$-group $G$ has the following three properties:
(i) $G\in\mathfrak{X}=\mathfrak{X}^{\prime}$, i.e. ${\tt GFlat}(kG)=\mathcal{F}(kG)$,
(ii) $G\in\mathfrak{Y}$, i.e. ${\tt GInj}(kG)=\mathcal{I}(kG)$ and
(iii) $kG$ is virtually Gorenstein.
Since $\mbox{sfli}\,kG$ is finite, all injective $kG$-modules have finite flat dimension. Hence, properties (i) and (ii) follow from the two general facts mentioned above; see also [39, Theorem 2.3 and Theorem 2.6]. As far as property (iii) is concerned, this is precisely [18, Theorem A]; see also [39, Lemma 3.2]. $\Box$
Let $\overline{\mathfrak{W}}$ be the smallest class of groups which contains $\mathfrak{W}$ and is ${\scriptstyle{{\bf LH}}}$-closed and $\Phi_{inj}$-closed. Groups in $\overline{\mathfrak{W}}$ admit a hierarchical description, as explained in §$\S 1$.IV. The following Corollary implies Theorem A stated in the Introduction; it is itself an immediate consequence of Theorem 4.2 and Proposition 4.3.
@@ 5. Groups satisfying Moore’s conjecture
Motivated by Serre’s theorem on torsion-free groups of finite virtual cohomological dimension [10, Chapter VIII, §$\S$3], Moore has conjectured that a module $M$ over the group algebra of a group $G$ is projective, whenever the restriction of $M$ to a subgroup $H\subseteq G$ of finite index that satisfies an additional group-theoretic condition is projective. More precisely, we say that the triple $(G,H,k)$, where $G$ is a group and $H\subseteq G$ is a subgroup of finite index, satisfies Moore’s condition if for any element $g\in G\setminus H$ one of the following two conditions hold:
@@
Proof. The first inclusion $\mathfrak{M}_{inj}\subseteq\mathfrak{M}_{flat}$ is a consequence of Lambek’s flatness criterion [31], as noted in [2, Proposition 2.6]. The inclusion $\mathfrak{M}_{flat}\subseteq\mathfrak{M}$ follows from the Benson-Goodearl theorem [8, Corollary 4.8], as in the proof of [2, Proposition 2.7]. $\Box$
@@
(iii) This follows from assertion (ii) above, in view of Lambek’s flatness criterion [31]. Indeed, the Pontryagin dual $DM$ of $M$ has injective dimension $\mbox{id}_{kG}DM=\mbox{fd}_{kG}M<\infty$, whereas the restricted $kH$-module $\mbox{res}_{H}^{G}DM=D\mbox{res}_{H}^{G}M$ is injective. We may then invoke (ii) and conclude that $DM\in{\tt Inj}(kG)$; hence, $M\in{\tt Flat}(kG)$.
(iv) Assume that $\mbox{fd}_{kG}M=n$ and let $N=\Omega^{n}M$ be the $n$-th syzygy in a projective resolution of $M$. Then, the $kG$-module $N$ is flat and its restriction $\mbox{res}_{H}^{G}N=\Omega^{n}\mbox{res}_{H}^{G}M$ is $kH$-projective. Invoking the Benson-Goodearl theorem [8, Corollary 4.8], we conclude that $N\in{\tt Proj}(kG)$ and hence $\mbox{pd}_{kG}M\leq n<\infty$. Then, the projectivity of $M$ is a consequence of (i) above. $\Box$
@@
(i) We shall first consider the case where $S$ is an ${\scriptstyle{{\bf H}}}\mathfrak{M}$-subgroup of $G$ and proceed by induction on the ordinal $\alpha$, for which $S\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{M}$, following the argument in the proof of [1, Proposition 4.2]. If $\alpha=0$, then $S\in\mathfrak{M}$ and hence $S$ satisfies Moore’s conjecture over $k$. The $kS$-module $N$ is therefore projective (as its restriction to $H\cap S$ is projective). For the inductive step, assume that $\alpha>0$ and the result is true for all ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{M}$-subgroups of $G$ and all $\beta<\alpha$. As we noted in §$\S$1.IV, there exists a short exact sequence of $kS$-modules
@@
We now assume that $S$ is an ${\scriptstyle{{\bf LH}}}\mathfrak{M}$-subgroup of $G$. If $S^{\prime}\subseteq S$ is a finitely generated subgroup, then there exists an ${\scriptstyle{{\bf H}}}\mathfrak{M}$-subgroup $S^{\prime\prime}\subseteq S$ with $S^{\prime}\subseteq S^{\prime\prime}$. Since the $kS^{\prime\prime}$-module $\mbox{res}_{S^{\prime\prime}}^{G}M$ is projective (in view of the above discussion, regarding ${\scriptstyle{{\bf H}}}\mathfrak{M}$-subgroups of $G$), this is also the case for its restriction $\mbox{res}_{S^{\prime}}^{G}M=\mbox{res}_{S^{\prime}}^{S}N$ to $S^{\prime}$; in particular, the $kS^{\prime}$-module $\mbox{res}_{S^{\prime}}^{S}N$ is flat. As $N$ is the filtered colimit of the $kS$-modules $\mbox{ind}_{S^{\prime}}^{S}\mbox{res}_{S^{\prime}}^{S}N\in{\tt Flat}(kS)$, where $S^{\prime}$ runs through all finitely generated subgroups of $S$, the $kS$-module $N$ is also flat. The projectivity of $N$ follows now from the Benson-Goodearl theorem [8, Corollary 4.8].
(ii) For any $\mathfrak{M}$-subgroup $F\subseteq S$, we consider the $kF$-module $\mbox{res}_{F}^{S}N=\mbox{res}_{F}^{G}M$ and note that its restriction $\mbox{res}_{H\cap F}^{F}N=\mbox{res}_{H\cap F}^{G}M=\mbox{res}_{H\cap S}^{H}\mbox{res}_{H}^{G}M$ to $H\cap F$ is projective (as a restriction of the projective $kH$-module $\mbox{res}_{H}^{G}M$). We also note that $(F,H\cap F,k)$ satisfies Moore’s condition, whereas the $\mathfrak{M}$-group $F$ satisfies Moore’s conjecture; it follows that $\mbox{res}_{F}^{S}N\in{\tt Proj}(kF)$. This holds for any $\mathfrak{M}$-subgroup $F$ of the $\Phi\mathfrak{M}$-group $S$ and hence $\mbox{pd}_{kS}N<\infty$. Then, Proposition 5.2(i) implies that $N\in{\tt Proj}(kS)$.
@@
(i) We first consider the case where $S$ is an ${\scriptstyle{{\bf H}}}\mathfrak{M}_{inj}$-subgroup of $G$ and proceed by induction on the ordinal $\alpha$, for which $S\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{M}_{inj}$, following the argument in the proof of [2, Theorem 2.5]. If $\alpha=0$, then $S\in\mathfrak{M}_{inj}$ and hence $S$ satisfies the injective version of Moore’s conjecture. The $kS$-module $N$ is therefore injective (as its restriction to $H\cap S$ is injective). For the inductive step, assume that $\alpha>0$ and the result is true for all ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{M}_{inj}$-subgroups of $G$ and all $\beta<\alpha$. As we noted in §$\S$1.IV, there exists a short exact sequence of $kS$-modules

### v2

@@ 4. Virtually Gorenstein group algebras
We fix a commutative ring $k$ and consider groups $G$ for which the group algebra $kG$ is virtually Gorenstein, i.e. for which ${\tt GProj}(kG)^{\perp}={{}^{\perp}}{\tt GInj}(kG)$. This homological condition on $kG$ has interesting consequences regarding the (homotopy) properties of Gorenstein modules. As an indication of these consequences, we note that if $kG$ is virtually Gorenstein, then: [^3] Properties (a), (b), (c) and (d) are actually valid over any (left) virtually Gorenstein ring $R$.
(a) (cf. [39, Proposition 3.16]) There is an equality ${\tt GProj}(kG)={\tt PGF}(kG)$; hence, [37, Theorem 4.4] implies that ${\tt GProj}(kG)\subseteq{\tt GFlat}(kG)$ and [37, Theorem 4.9] implies that the cotorsion pair $\left({\tt GProj}(kG),{\tt GProj}(kG)^{\perp}\right)$ is complete.
(b) (cf. [12, Proposition 2.6]) Let $M$ be a $kG$-module. Any resolution of $M$ which is obtained by iterating special Gorenstein projective precovers remains acyclic by applying the functor $\mbox{Hom}_{kG}(\_\!\_,N)$ for any Gorenstein injective $kG$-module $N$. Dually, any coresolution of $M$ which is obtained by iterating special Gorenstein injective preenvelopes remains acyclic by applying the functor $\mbox{Hom}_{kG}(L,\_\!\_)$ for any Gorenstein projective $kG$-module $L$.
(c) (cf. [12, Proposition 2.2]) An acyclic complex of $kG$-modules remains acyclic by applying the functor $\mbox{Hom}_{kG}(\_\!\_,N)$ for any Gorenstein injective $kG$-module $N$ if and only if it remains acyclic by applying the functor $\mbox{Hom}_{kG}(L,\_\!\_)$ for any Gorenstein projective $kG$-module $L$.
(d) (cf. [39, Example 3.18(2)]) Let $\mathcal{C}$ be the class of those acyclic complexes of $kG$-modules that satisfy the equivalent conditions described in (c) above. Then, the relative derived category, which is obtained from the homotopy category of $kG$ by inverting those chain maps whose mapping cone is a $\mathcal{C}$-complex, is triangulated equivalent to certain homotopy categories of Gorenstein projective (resp. Gorenstein injective) $kG$-modules.
We consider the classes $\mathfrak{X}$ and $\mathfrak{Y}$ defined in Section 2 and let $\mathfrak{Z}=\mathfrak{Z}(k)$ be the class consisting of those $\mathfrak{Y}$-groups $G$, for which $kG$ is virtually Gorenstein.
@@
Proof. Let $G$ be a $\mathfrak{Z}$-group and consider a subgroup $H\subseteq G$. Then, $G\in\mathfrak{Y}$ and $kG$ is virtually Gorenstein. Since $\mathfrak{Y}$ is subgroup-closed (cf. Theorem 2.1), it follows that $H\in\mathfrak{Y}$. It remains to show that $kH$ is virtually Gorenstein.
Consider a $kH$-module $M\in{\tt GProj}(kH)^{\perp}$. Since $H\in\mathfrak{Y}\subseteq\mathfrak{X}$ (cf. Theorem 2.1(i)), restriction maps ${\tt GProj}(kG)$ into ${\tt GProj}(kH)$; this is precisely property (4) in Section 2. It follows that the coinduced $kG$-module $\mbox{coind}_{H}^{G}M$ is contained in ${\tt GProj}(kG)^{\perp}$, i.e. in ${{}^{\perp}}{\tt GInj}(kG)$. Invoking Corollary 2.2(ii), we may now conclude that $\mbox{res}_{H}^{G}\mbox{coind}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$. Since $M$ is a direct summand of $\mbox{res}_{H}^{G}\mbox{coind}_{H}^{G}M$, it follows that $M\in{{}^{\perp}}{\tt GInj}(kH)$. Hence, we have shown that ${\tt GProj}(kH)^{\perp}\subseteq{{}^{\perp}}{\tt GInj}(kH)$.
Conversely, assume that $M\in{{}^{\perp}}{\tt GInj}(kH)$. Since $H$ is a $\mathfrak{Y}$-group, restriction maps ${\tt GInj}(kG)$ into ${\tt GInj}(kH)$; this is precisely property (4’) in Section 2. It follows that the induced $kG$-module $\mbox{ind}_{H}^{G}M$ is contained in ${{}^{\perp}}{\tt GInj}(kG)$, i.e. in ${\tt GProj}(kG)^{\perp}$. Corollary 2.2(i) then implies that $\mbox{res}_{H}^{G}\mbox{ind}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$. Since $M$ is a direct summand of $\mbox{res}_{H}^{G}\mbox{ind}_{H}^{G}M$, it follows that $M\in{\tt GProj}(kH)^{\perp}$. We have thus shown that ${{}^{\perp}}{\tt GInj}(kH)\subseteq{\tt GProj}(kH)^{\perp}$. $\Box$
@@
Proof. (i) We have to show that ${\scriptstyle{{\bf LH}}}\mathfrak{Z}=\mathfrak{Z}$, i.e. that ${\scriptstyle{{\bf LH}}}\mathfrak{Z}\subseteq\mathfrak{Z}$. To that end, we consider a group $G\in{\scriptstyle{{\bf LH}}}\mathfrak{Z}$ and aim at proving that $G\in\mathfrak{Z}$. We note that ${\scriptstyle{{\bf LH}}}\mathfrak{Z}\subseteq{\scriptstyle{{\bf LH}}}\mathfrak{Y}=\mathfrak{Y}$ and hence $G\in\mathfrak{Y}$. It remains to show that $kG$ is virtually Gorenstein. For any $\mathfrak{Z}$-subgroup $H\subseteq G$ we denote by $\left(\mbox{res}_{H}^{G}\right)^{-1}{\tt GProj}(kH)^{\perp}$ the class of those $kG$-modules $M$ for which $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ and apply Corollary 3.2 (with $\mathfrak{X}_{0}=\mathfrak{Z}\subseteq\mathfrak{Y}\subseteq\mathfrak{X}$ therein), in order to conclude that
@@
Since ${\tt GProj}(kH)^{\perp}={{}^{\perp}}{\tt GInj}(kH)$ for any $\mathfrak{Z}$-subgroup $H\subseteq G$, it follows that ${\tt GProj}(kG)^{\perp}={{}^{\perp}}{\tt GInj}(kG)$. The group algebra $kG$ is therefore virtually Gorenstein and hence $G\in\mathfrak{Z}$.
(ii) In order to show that $\Phi_{inj}\mathfrak{Z}=\mathfrak{Z}$, i.e. that $\Phi_{inj}\mathfrak{Z}\subseteq\mathfrak{Z}$, we consider a $\Phi_{inj}\mathfrak{Z}$-group $G$ and aim at proving that $G\in\mathfrak{Z}$. First of all, we note that $\Phi_{inj}\mathfrak{Z}\subseteq\Phi_{inj}\mathfrak{Y}=\mathfrak{Y}$, so that $G\in\mathfrak{Y}$. Since $G$ is contained in $\Phi_{inj}\mathfrak{Z}\subseteq\Phi_{flat}\mathfrak{Z}$ (cf. [23, Lemma 1.5]), we can apply Corollaries 3.2 and 3.5 in exactly the same way as in (i) above and show that ${\tt GProj}(kG)^{\perp}={{}^{\perp}}{\tt GInj}(kG)$, i.e. that the group algebra $kG$ is virtually Gorenstein. $\Box$
Let $\mathfrak{W}=\mathfrak{W}(k)$ be the class consisting of those groups $G$, for which $kG$ is weakly Gorenstein regular; in other words, $G\in\mathfrak{W}$ if and only if $\mbox{sfli}\,kG<\infty$. This class contains all groups $G$, for which $kG$ is Gorenstein regular. We note that $\mathfrak{W}$ is subgroup-closed; indeed, for any subgroup $H$ of a group $G$ there is an inequality $\mbox{sfli}\,kH\leq\mbox{sfli}\,kG$. Hence, $\mathfrak{W}$ is non-empty if and only if it contains the trivial group, i.e. if and only if $k$ is itself weakly Gorenstein regular. In that case, $\mathfrak{W}$ is known to contain the class $\mathfrak{F}$ of finite groups; see, for example, [35].
@@
Proof. We have to show that any $\mathfrak{W}$-group $G$ has the following two properties: (i) $G\in\mathfrak{Y}$, i.e. ${\tt GInj}(kG)=\mathcal{I}(kG)$ and (ii) $kG$ is virtually Gorenstein.
Since $\mbox{sfli}\,kG<\infty$, all injective $kG$-modules have finite flat dimension. On the other hand, if $I^{*}$ is any acyclic complex of injective $kG$-modules and $M$ is any $kG$-module of finite flat dimension, then the complex of abelian groups $\mbox{Hom}_{kG}(M,I^{*})$ is acyclic; cf. [36, Corollary 5.9]. This shows that $G$ satisfies property (i) above; see also [41, Theorem 2.6]. As far as property (ii) is concerned, this is precisely [19, Theorem A]; see also [41, Lemma 3.2]. $\Box$
Let $\overline{\mathfrak{W}}$ be the smallest class of groups which contains $\mathfrak{W}$ and is ${\scriptstyle{{\bf LH}}}$-closed and $\Phi_{inj}$-closed. Groups in $\overline{\mathfrak{W}}$ admit a hierarchical description, as explained in §$\S 1$.IV. The following Corollary implies Theorem A in the Introduction; it is itself an immediate consequence of Theorem 4.2 and Proposition 4.3.
@@ 5. Groups satisfying Moore’s conjecture
Motivated by Serre’s theorem on torsion-free groups of finite virtual cohomological dimension [11, Chapter VIII, §$\S$3], Moore has conjectured that a module $M$ over the group algebra of a group $G$ is projective, whenever the restriction of $M$ to a subgroup $H\subseteq G$ of finite index that satisfies an additional group-theoretic condition is projective. More precisely, we say that the triple $(G,H,k)$, where $G$ is a group and $H\subseteq G$ is a subgroup of finite index, satisfies Moore’s condition if for any element $g\in G\setminus H$ one of the following two conditions hold:
@@
Proof. The first inclusion $\mathfrak{M}_{inj}\subseteq\mathfrak{M}_{flat}$ is a consequence of Lambek’s flatness criterion [33], as noted in [3, Proposition 2.6]. The inclusion $\mathfrak{M}_{flat}\subseteq\mathfrak{M}$ follows from the Benson-Goodearl theorem [9, Corollary 4.8], as in the proof of [3, Proposition 2.7]. $\Box$
@@
(iii) This follows from assertion (ii) above, in view of Lambek’s flatness criterion [33]. Indeed, the Pontryagin dual $DM$ of $M$ has injective dimension $\mbox{id}_{kG}DM=\mbox{fd}_{kG}M<\infty$, whereas the restricted $kH$-module $\mbox{res}_{H}^{G}DM=D\mbox{res}_{H}^{G}M$ is injective. We may then invoke (ii) and conclude that $DM\in{\tt Inj}(kG)$, so that $M\in{\tt Flat}(kG)$.
(iv) Assume that $\mbox{fd}_{kG}M=n$ and let $N=\Omega^{n}M$ be the $n$-th syzygy in a projective resolution of $M$. Then, the $kG$-module $N$ is flat and its restriction $\mbox{res}_{H}^{G}N=\Omega^{n}\mbox{res}_{H}^{G}M$ is $kH$-projective. Invoking the Benson-Goodearl theorem [9, Corollary 4.8], we conclude that $N\in{\tt Proj}(kG)$ and hence $\mbox{pd}_{kG}M\leq n<\infty$. Then, the projectivity of $M$ is a consequence of (i) above. $\Box$
@@
(i) We shall first consider the case where $S$ is an ${\scriptstyle{{\bf H}}}\mathfrak{M}$-subgroup of $G$ and proceed by induction on the ordinal $\alpha$, for which $S\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{M}$, following the argument in the proof of [2, Proposition 4.2]. If $\alpha=0$, then $S\in\mathfrak{M}$ and hence $S$ satisfies Moore’s conjecture over $k$. The $kS$-module $N$ is therefore projective (as its restriction to $H\cap S$ is projective). For the inductive step, assume that $\alpha>0$ and the result is true for all ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{M}$-subgroups of $G$ and all $\beta<\alpha$. As we noted in §$\S$1.IV, there exists a short exact sequence of $kS$-modules
@@
We now assume that $S$ is an ${\scriptstyle{{\bf LH}}}\mathfrak{M}$-subgroup of $G$. If $S^{\prime}\subseteq S$ is a finitely generated subgroup, then there exists an ${\scriptstyle{{\bf H}}}\mathfrak{M}$-subgroup $S^{\prime\prime}\subseteq S$ with $S^{\prime}\subseteq S^{\prime\prime}$. Since the $kS^{\prime\prime}$-module $\mbox{res}_{S^{\prime\prime}}^{G}M$ is projective (in view of the above discussion, regarding ${\scriptstyle{{\bf H}}}\mathfrak{M}$-subgroups of $G$), this is also the case for its restriction $\mbox{res}_{S^{\prime}}^{G}M=\mbox{res}_{S^{\prime}}^{S}N$ to $S^{\prime}$; in particular, the $kS^{\prime}$-module $\mbox{res}_{S^{\prime}}^{S}N$ is flat. As $N$ is the filtered colimit of the $kS$-modules $\mbox{ind}_{S^{\prime}}^{S}\mbox{res}_{S^{\prime}}^{S}N\in{\tt Flat}(kS)$, where $S^{\prime}$ runs through all finitely generated subgroups of $S$, the $kS$-module $N$ is also flat. The projectivity of $N$ follows now from the Benson-Goodearl theorem [9, Corollary 4.8].
(ii) For any $\mathfrak{M}$-subgroup $F\subseteq S$, we consider the $kF$-module $\mbox{res}_{F}^{S}N=\mbox{res}_{F}^{G}M$ and note that its restriction $\mbox{res}_{H\cap F}^{S}N=\mbox{res}_{H\cap F}^{G}M=\mbox{res}_{H\cap F}^{H}\mbox{res}_{H}^{G}M$ to $H\cap F$ is projective (as a restriction of the projective $kH$-module $\mbox{res}_{H}^{G}M$). We also note that $(F,H\cap F,k)$ satisfies Moore’s condition, whereas the $\mathfrak{M}$-group $F$ satisfies Moore’s conjecture; it follows that $\mbox{res}_{F}^{S}N\in{\tt Proj}(kF)$. This holds for any $\mathfrak{M}$-subgroup $F$ of the $\Phi\mathfrak{M}$-group $S$ and hence $\mbox{pd}_{kS}N<\infty$. Then, Proposition 5.2(i) implies that $N\in{\tt Proj}(kS)$.
@@
(i) We first consider the case where $S$ is an ${\scriptstyle{{\bf H}}}\mathfrak{M}_{inj}$-subgroup of $G$ and proceed by induction on the ordinal $\alpha$, for which $S\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{M}_{inj}$, following the argument in the proof of [3, Theorem 2.5]. If $\alpha=0$, then $S\in\mathfrak{M}_{inj}$ and hence $S$ satisfies the injective version of Moore’s conjecture. The $kS$-module $N$ is therefore injective (as its restriction to $H\cap S$ is injective). For the inductive step, assume that $\alpha>0$ and the result is true for all ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{M}_{inj}$-subgroups of $G$ and all $\beta<\alpha$. As we noted in §$\S$1.IV, there exists a short exact sequence of $kS$-modules